I need help on solving equations with variables on both sides like:
x-3=6x+7

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

x-3=6x+7
Subtract "x" from both sides to get:
-3=5x+7
Subtract 7 from both sides to get:
-10=5x
Divide both sides by 5 to get:
x=-2
Cheers,
Stan H.

Answer by checkley71(8403)

X-3=6X+7
YOU NEED TO GET ALL VARIABLES ON THE SAME SIDE OF THE EQUATION AND IF POSSIBLE COMBINE THEM AND GET THE NUMBERS ON THE OPPOSITE SIDE OF THE EQUATION
X-6X=7+3
-5X=10 NO DIVIDE BOTH SIDES BY -5
X=10/-5
X=-2
PROOF
-2-3=6*-2+7
-5=-12+7
-5=-5